Use the Laplace transform to solve the given initial-value problem. y + y = V2 sin it, y(0) = 13, y(0) = 0 y(t) =
Answer #1

Date: il Page No – yll ty = 32 sin et o Apply Laplace transform on both sides L29} +L{y} = JR LS sio Jet? [S2 Ycs) - 5.90–90> 1332 +265 + = CASTB2653+13 + (CS+D 28+2) $ 1383 +265 +1 = CAS3 +AS +38+B)+(Cs>t2 cs tos +2D) 138 +265+ (Atc213 +CB+D2S2+(A1.y(t): 26 Id LIII 2 13eds Ladin 5242 J (52+2 (3+1 800 = 26 cosat - cost-13 (sint)+ (sint) 90 = 2500szt -12sipt.